OpenStudy (anonymous):

Please explain this, step by step preferably. Mark and Ely can paint a room in 4 hours. Pepe can paint the same room in 4 hours. How long will it take to paint if the 3 of them work together and Mark words as fast as Ely?

OpenStudy (anonymous):

could it be like this: (4+4)/3= 2/2/3 hours

OpenStudy (anonymous):

in 1 hour mark and ely can paint 1/4 of room in 1 hour pepe can paint 1/4 of room in 1 hour they can together paint = 1/4 + 1/4 = 2/4 = 1/2 of the room So to paint the entire room they need 2 hours

OpenStudy (anonymous):

hello